# Benchmarking pig detection and tracking under diverse and challenging conditions

### v1

@@ 2.3 Dataset splits
Images from one barn (N = 250) were selected as the test set and show an environment that is not present in the training (N = 2431) and validation data (N = 250). This independence of our test set was chosen to challenge the robustness of models, as well as simulate practical conditions where models have to perform in previously unseen environments. Apart from randomly sampled images (N = 130, 13 night images), the test dataset contains a substantial portion of challenging images (N = 120, 27 night images). We chose to include a large proportion of challenging images to enable a better differentiation between the performance of different methods. Tab. 1 includes the number of images, challenging images, night images, and annotated bounding boxes for each dataset split. A depiction of the pen environment of the test set can be found in Fig. 8 (pen setting H6).
@@ 3.1 Annotation
The videos in our PigTrack dataset originate from the same video footage that has been used to obtain images for PigDetect (see Sec. 2.1). To obtain videos for annotation, we first sighted the footage to specifically identify segments that featured challenging tracking conditions (no formal criterion). To avoid potential ambiguities during annotation, it was further ensured that no pig leaves the scene and re-enters afterwards. Pigs leaving without returning or new pigs entering throughout the video was considered acceptable. The selected video segments were input to a basic tracking method to generate preliminary annotations. Specifically, we used YOLOX [53] to generate frame-wise bounding boxes and linked them using a simple motion-based matching algorithm (C-BIoU, [55]). The resulting preliminary annotations were loaded into CVAT [56], an image and video annotation tool, so that only the errors had to be manually corrected. This protocol greatly improved annotation efficiency since correctly predicted bounding boxes and (sub-)tracks could be left unchanged, allowing annotators to focus on refining only the challenging parts. Tracks were annotated by one person and independently reviewed by a second person, with corrections applied if necessary. Five of the authors were involved in this procedure (J. H., C. P., P. S., E. C. and A. M. Y.). In total, 80 videos were densely annotated (every frame), spanning 41.06 min of material. As with PigDetect, some particularly complex cases could not be unambiguously annotated and were therefore not included in the dataset.
@@ 3.2 Characteristics
Dataset T (min) #Vids #Night #Frames #IDs #BBoxes #Pens Type
Bergamini et al. [10] 60.00 12 0 21 6002160021\,60021 600 96 12 6461264612\,64612 646 1 sparse
Melfsen et al. [30] 14.00 14 3 25 2002520025\,20025 200 154 9240 1 sparse
Jaoukaew et al. [40] 9.33 4 0 19 3391933919\,33919 339 41 554 1 sparse
Yu et al. [42] 2.50 1 0 1260 23 28 8142881428\,81428 814 1 dense
PigTrack (our) 41.06 80 26 26 4872648726\,48726 487 1574 517 961517961517\,961517 961 9 dense
train 15.12 35 13 9959 698 195 745195745195\,745195 745 9 dense
val 12.06 20 3 7589 383 141 289141289141\,289141 289 7 dense
test 13.88 25 10 8939 493 180 927180927180\,927180 927 9 dense
Tab. 2 summarizes the characteristics of PigTrack in comparison to existing pig tracking datasets. PigTrack contains substantially more videos, night recordings, unique pig IDs, and annotated bounding boxes than any previously published dataset. While some earlier datasets have longer durations [10] or a comparable number of frames [10, 30, 40], they are only sparsely annotated, resulting in far fewer bounding box labels. Specifically, Bergamini et al. [10] annotated frames irregularly based on when behaviors change, Melfsen et al. [30] annotated only one frame per second, and Jaoukaew et al. [40] only every 500th frame. Standard benchmarks for human multi-object tracking are usually densely labeled [57, 49]. While sparsely labeled bounding boxes can be interpolated for training or evaluation purposes, the quality of tracks will inevitably decrease compared to dense annotations. Moreover, most existing pig tracking datasets [10, 30, 40, 42] are limited to a single pen environment. A comprehensive depiction of the visual diversity of the pen settings of previously published datasets (Fig. 9) and PigTrack (Fig. 10) can be found in B. For more detailed characteristics for every video included in PigTrack, we refer to Tab. 10.
The width and height of most videos (Inno Pig, FriSch, Old Breed New House) is $1280\times 800$ , while the videos from one source (TiPP) have a resolution of $1920\times 1080$ pixels. The majority of videos (N = 62) has 10 frames per second (FPS), while a smaller number operate at either 5 FPS (N = 2) or 15 FPS (N = 16). The dataset features nine distinct pen environments and includes 26 night videos. All pigs visible in a video were annotated. On rare occasions, this included animals in neighboring pens. When pigs in neighboring pens could not be clearly annotated, the corresponding regions were masked by setting their pixel values to zero. The breed of pigs as well as the pen and camera characteristics have already been described in Sec. 2.2.
@@ 4.1 Evaluated object detection models
We benchmarked test performance using the two real-time object detection models YOLOX [53] and YOLOv8 [58]. Furthermore, we benchmarked the performance of DINO [59] with collaborative hybrid assignments training [46] (Co-DINO), a model focused on detection quality rather than inference speed. These three models were selected for benchmarking because they are widely adopted, have a strong performance, and are supported by the software framework used in our study (see Sec. 4.1.3). Co-DINO achieves state-of-the-art results on the COCO dataset [45], but has not yet been employed for pig detection.
4.1.1 YOLO family
@@
In the original YOLO model proposed by Redmon et al. [60], an input image is processed by a convolutional backbone to produce a rich visual feature map of size $7\times 7\times 1024$. Intuitively, this can be thought of as corresponding to a $7\times 7$ grid over the input image, while the feature vector of size 1024 contains all relevant information for objects at the corresponding grid location. The feature vector is subsequently processed by a smaller head network to predict class probabilities and bounding box coordinates relative to the grid position. Compared to more complex proposal-based paradigms, as used by the R-CNN family [61, 62, 63, 14], the YOLO model is computationally efficient, enabling real-time object detection.
@@
Successor models like YOLOX [53] and YOLOv8 [58] made several improvements over the original model. For example, both models adopt separate head branches for classification and bounding box regression. This helps prevent training conflicts between these semantically distinct tasks by allowing each branch to learn its own features. Notable performance increases have also been achieved by incorporating strong data augmentations. For example, multiple images are overlayed (Mixup, [64]) or combined (Mosaic, [65]) to create more challenging prediction scenarios. Other improvements include loss functions that empirically lead to an increased performance [66, 67] or better label assignment strategies between predicted and ground-truth bounding boxes for more effective training [53].
4.1.2 Co-DINO
@@
Co-DINO is based on the seminal DETR [68] architecture for object detection. In the DETR architecture, a convolutional backbone first extracts features from the input image, which are subsequently refined by a transformer [69] encoder. Then, a set of learned queries extracts information from these features via cross-attention in a transformer decoder architecture. The transformer decoder also includes self-attention between queries, which is crucial to prevent multiple queries from attending to the same object, thereby avoiding duplicate bounding boxes. Finally, the enriched query features are used for classification and bounding box regression. For supervision, a one-to-one matching needs to be established between queries and ground-truth objects. This is done by choosing the matching that minimizes the loss between the predicted and ground-truth bounding boxes and classes using the Hungarian algorithm [70].
@@
The original DETR model suffered from convergence problems and performed worse than fully convolutional object detection approaches. Follow-up works [71, 72, 73] proposed several improvements, which have been combined and extended in the DINO model [59]. For example, DINO parameterizes queries as dynamic anchor boxes [73], treating the bounding box regression task as an iterative refinement of these anchors across multiple transformer decoder layers. This spatial informativeness of queries promotes convergence and boosts performance.
@@
Co-DINO [46] further improves upon the training recipe of DINO. In DINO (and other DETR frameworks), the number of positive queries is limited by the number of objects in an image. This leads to a sparse supervision of the backbone, which is detrimental for feature learning. Co-DINO tackles this by including auxiliary object detection heads such as Faster-RCNN [63] for supervision. The small number of positive queries also leads to inefficient cross-attention learning. This is tackled by generating customized positive queries based on bounding boxes obtained from the auxiliary heads. Together, these advancements establish Co-DINO as the state of the art in object detection.
@@ 4.1.3 Implementation details
Table 3: Model and training specifications for pig detection. Pretrained Model indicates the dataset on which the model used for weight initialization was trained. The learning rate schedule for Co-DINO decays by a factor of 0.1 at the step. The training time for YOLOX and YOLOv8 is reported for the largest model variant.
Model Backbone Pretrained Model Batch Size Epochs LR Schedule LR (max/min) Train Time (h)
YOLOX DarkNet53[74] COCO[45] 6 115 Cosine $\frac{6}{64}\times 10^{-2}\;\big{/}\;\frac{30}{64}\times 10^{-4}$ 15.1
YOLOv8 CSP-Darknet[75] COCO 6 110 Linear $\frac{6}{64}\times 10^{-2}\;\big{/}\;\frac{6}{64}\times 10^{-4}$ 9.7
Co-DINO Swin-L[76] COCO + Objects365[77] 4 100 Step (at 90) $\frac{1}{4}\times 10^{-4}\;\big{/}\;\frac{1}{4}\times 10^{-5}$ 18.5
All models were trained and validated according to the dataset splits defined in Sec. 2.3. The best-performing model in terms of mean Average Precision (see Sec. 4.3) on the validation set was then used to obtain test performance. YOLOX and YOLOv8 were trained on a single Nvidia A100 GPU using the MMYolo toolbox [78], while Co-DINO was trained on four A100 GPUs using the MMDetection toolbox [79]. All models use an image resolution of $960\times 960$ pixels. The basic training configurations are listed in Tab. 3. Other specifications such as learning rate warmup schedules or data augmentations are in line with the original implementations.
The speed of the trained models was measured on an Nvidia A100 GPU using the inference functionality of the MMDetection toolbox. We ran the models on a set of 1000 randomly selected images, repeating the process five times to mitigate variability in performance measurements. The overall median speed per image was computed from these 5000 inferences to obtain a robust assessment. Images were resized in advance to a resolution of $960\times 960$ pixels to avoid additional overhead. Data loading from disc was not considered in the speed measurements, as it can be avoided in practical applications.
@@
The model used for training is YOLOX [53] using a checkpoint trained on COCO as initialization . A constant learning rate of $3/64\times 10^{-2}$ with a batch size of 3 was used in all 14 variants to ensure comparability. The model was trained with strong data augmentations throughout (Mosaic and Mixup). Otherwise, the training settings were identical to the ones that were used to train the “YOLOX-L” model in the original YOLOX work [53]. Training was terminated when the mean Average Precision (see Sec. 4.3) on the test set did not increase for 20 consecutive epochs. Performance metrics are reported for the checkpoint with the highest mean Average Precision score.
@@ 4.3 Evaluation metrics
Model performance was evaluated with the commonly used Average Precision metric [80]. To construct this metric, a one-to-one matching between predicted and ground-truth bounding boxes is established by minimizing the IoU across all pairs. This is done using the Hungarian algorithm [70]. The algorithm ensures that each predicted bounding box can be matched to at most one ground-truth bounding box and vice versa. The resulting matches are subsequently filtered according to a minimum required IoU of $\alpha$. This prevents matches of predictions and ground truths that have little or no overlap. Next, all $N_{\text{pred}}$ predicted bounding boxes in the dataset that is evaluated are ordered by confidence in descending order. Then, precision and recall based on the $n$ most confident predictions and a specific $\alpha$ value can be calculated as:
@@
[…]
Following the literature on object detection, we report two metrics in this work: $\text{AP}_{0.5}$ and mean Average Precision (AP). $\text{AP}_{0.5}$ is simply $\text{AP}_{\alpha}$ computed with a minimum required IoU for a match of $\alpha=0.5$. Since the required IoU is relatively low, this metric measures a model’s ability to roughly detect objects without taking into account localization accuracy. In contast, AP is computed by averaging over multiple IoU thresholds:

### v2

@@ 2.3 Dataset splits
Images from one barn (N = 250) were selected as the test set and show an environment that is not present in the training (N = 2431) and validation data (N = 250). This independence of our test set was chosen to challenge the robustness of models, as well as simulate practical conditions where models have to perform in previously unseen environments. Apart from randomly sampled images (N = 130, 13 night images), the test set contains a substantial portion of challenging images (N = 120, 27 night images). We chose to include a large proportion of challenging images to enable a better differentiation between the performance of different methods. Tab. 1 includes the number of images, challenging images, night images, and annotated bounding boxes for each dataset split. A depiction of the pen environment of the test set can be found in Fig. 8 (pen setting H6).
@@ 3.1 Annotation
The videos in our PigTrack dataset originate from the same video footage that has been used to obtain images for PigDetect (see Sec. 2.1). At the time of annotation, we were not aware of publicly available video datasets with a license that permits republication of modified versions, which is why we did not enrich the diversity of our dataset with third-party videos. To obtain videos for annotation, we first sighted the footage to specifically identify segments that featured challenging tracking conditions (no formal criterion). This included scenarios with rapid, non-linear movements, poor visibility, and occlusions. In case of temporary occlusions, it was ensured that all pigs could still be unambiguously identified throughout the video. For example, this is the case if only one pig is occluded at a time, if a pig reappears in the same location, or if it follows a clear trajectory during the occlusion. To avoid further ambiguities, it was also ensured that no pig leaves the scene and re-enters afterwards. Pigs leaving without returning or new pigs entering throughout the video was considered acceptable.
The selected video segments were input to a basic tracking method to generate preliminary annotations. Specifically, we used YOLOX [ge_yolox_2021] to generate frame-wise bounding boxes and linked them using a simple motion-based matching algorithm (C-BIoU, [yang_hard_2023]). The resulting preliminary annotations were loaded into CVAT [cvat], an image and video annotation tool, so that only the errors had to be manually corrected. This protocol greatly improved annotation efficiency since correctly predicted bounding boxes and (sub-)tracks could be left unchanged, allowing annotators to focus on refining only the challenging parts. CVAT furthermore allows hiding all tracks except one, such that the videos could be annotated one pig at a time, which prevents sensory overload. Tracks were annotated by one person and independently reviewed by a second person, with corrections applied if necessary. Five of the authors were involved in this procedure (J. H., C. P., P. S., E. C. and A. M. Y.). In total, 80 videos were densely annotated (every frame), spanning 41.06 min of material. As with PigDetect, some particularly complex cases could not be unambiguously annotated and were therefore not included in the dataset.
@@ 3.2 Characteristics
Dataset T (min) Videos Night Frames Tracks BBoxes Pens Type
bergamini_extracting_2021 60.00 12 0 21 60021\,600 96 12 64612\,646 1 sparse
melfsen2023describing 14.00 14 3 25 20025\,200 154 9240 1 sparse
jaoukaew2024robust 9.33 4 0 19 33919\,339 41 554 1 sparse
yu2025fto 2.50 1 0 1260 23 28 81428\,814 1 dense
PigTrack (our) 41.06 80 26 26 48726\,487 1574 517 961517\,961 9 dense
train 15.12 35 13 9959 698 195 745195\,745 9 dense
val 12.06 20 3 7589 383 141 289141\,289 7 dense
test 13.88 25 10 8939 493 180 927180\,927 9 dense
Tab. 2 summarizes the characteristics of PigTrack in comparison to existing pig tracking datasets. PigTrack contains substantially more videos, night recordings, tracks, and annotated bounding boxes than any previously published dataset. While some earlier datasets have longer durations [bergamini_extracting_2021] or a comparable number of frames [bergamini_extracting_2021, melfsen2023describing, jaoukaew2024robust], they are only sparsely annotated, resulting in far fewer bounding box labels. Specifically, bergamini_extracting_2021 annotated frames irregularly based on when behaviors change, melfsen2023describing annotated only one frame per second, and jaoukaew2024robust only every 500th frame. Standard benchmarks for human multi-object tracking are usually densely labeled [milan2016mot16, sun2022dancetrack]. While sparsely labeled bounding boxes can be interpolated for training or evaluation purposes, the quality of tracks will inevitably decrease compared to dense annotations. Moreover, existing pig tracking datasets [bergamini_extracting_2021, melfsen2023describing, jaoukaew2024robust, yu2025fto] are limited to a single pen environment. A comprehensive depiction of the visual diversity of the pen settings of previously published datasets (Fig. 11) and PigTrack (Fig. 12) can be found in B. For more detailed characteristics for every video included in PigTrack, we refer to Tab. 11.
The width and height of most videos (Inno Pig, FriSch, Old Breed New House) is $1280\times 800$ , while the videos from one source (TiPP) have a resolution of $1920\times 1080$ pixels. The majority of videos (N = 62) has 10 frames per second (FPS), while a smaller number operate at either 5 FPS (N = 2) or 15 FPS (N = 16). The dataset features nine distinct pen environments and includes 26 night videos, which were captured using infrared illumination. All pigs visible in a video were annotated. On rare occasions, this included animals in neighboring pens. When pigs in neighboring pens could not be clearly annotated, the corresponding regions were masked by setting their pixel values to zero. If pigs are temporarily invisible in a video sequence (e. g., due to occlusion), only those frames in which they are visible were annotated, following established multi-object tracking annotation protocols [milan2016mot16, sun2022dancetrack]. The breed of pigs as well as the pen and camera characteristics have already been described in Sec. 2.2.
@@ 4.1 Evaluated object detection models
We benchmarked test performance using the three real-time object detection models YOLOX [ge_yolox_2021], YOLOv8 [Jocher_Ultralytics_YOLO_2023] and RT-DETR [zhao2024detrs]. Furthermore, we benchmarked the performance of DINO [zhang2022dino] with collaborative hybrid assignments training [zong2023detrs] (Co-DINO), a model focused on detection quality rather than inference speed. These four models were selected for benchmarking because they are widely adopted and have a strong performance. Furthermore, YOLOX, YOLOv8, and Co-DINO are all supported within the same software framework, which we used as the foundation of our code base (see Sec. 4.1.3). Co-DINO achieves state-of-the-art results on the COCO dataset [lin2014microsoft], but has not yet been employed for pig detection.
4.1.1 YOLOX and YOLO-v8
@@
In the original YOLO model proposed by redmon2016you, an input image is processed by a convolutional backbone to produce a rich visual feature map of size $7\times 7\times 1024$. Intuitively, this can be thought of as corresponding to a $7\times 7$ grid over the input image, while the feature vector of size 1024 contains all relevant information for objects at the corresponding grid location. The feature vector is subsequently processed by a smaller head network to predict class probabilities and bounding box coordinates relative to the grid position. Compared to more complex proposal-based paradigms, as used by the R-CNN family [girshick2014rich, girshick2015fast, ren2015faster, he2017mask], the YOLO model is computationally efficient, enabling real-time object detection.
@@
Successor models like YOLOX [ge_yolox_2021] and YOLOv8 [Jocher_Ultralytics_YOLO_2023] made several improvements over the original model. For example, both models adopt separate head branches for classification and bounding box regression. This helps prevent training conflicts between these semantically distinct tasks by allowing each branch to learn its own features. Notable performance increases have also been achieved by developing more elaborate data augmentations. For example, multiple images are overlayed (Mixup, [zhang2019bag]) or combined (Mosaic, [bochkovskiy2020yolov4]) to create more challenging prediction scenarios. Other improvements include loss functions that empirically lead to an increased performance [li2020generalized, zheng2020distance] or better label assignment strategies between predicted and ground-truth bounding boxes for more effective training [ge_yolox_2021].
4.1.2 Co-DINO and RT-DETR
@@
Both Co-DINO and RT-DETR are based on the seminal DETR [carion2020end] architecture for object detection. In the DETR architecture, a convolutional backbone first extracts features from the input image, which are subsequently refined by a transformer [vaswani2017attention] encoder. Then, a set of learned queries extracts information from these features via cross-attention in a transformer decoder architecture. The transformer decoder also includes self-attention between queries, which is crucial to prevent multiple queries from attending to the same object, thereby avoiding duplicate bounding boxes. Finally, the enriched query features are used for classification and bounding box regression. For supervision, a one-to-one matching needs to be established between queries and ground-truth objects. This is done by choosing the matching that minimizes the loss between the predicted and ground-truth bounding boxes and classes using the Hungarian algorithm [kuhn1955hungarian].
@@
The original DETR model suffered from convergence problems and performed worse than fully convolutional object detection approaches. Follow-up works [zhu2020deformable, li2022dn, liu2022dab] proposed several improvements, which have been combined and extended in the DINO model [zhang2022dino]. For example, DINO parameterizes queries as dynamic anchor boxes [liu2022dab], treating the bounding box regression task as an iterative refinement of these anchors across multiple transformer decoder layers. This spatial informativeness of queries promotes convergence and boosts performance.
@@
Co-DINO [zong2023detrs] further improves upon the training recipe of DINO. In DINO (and other DETR frameworks), the number of positive queries is limited by the number of objects in an image. This leads to a sparse supervision of the backbone, which is detrimental for feature learning. Co-DINO tackles this by including auxiliary object detection heads such as Faster-RCNN [ren2015faster] for supervision. The small number of positive queries also leads to inefficient cross-attention learning. This is tackled by generating customized positive queries based on bounding boxes obtained from the auxiliary heads. Together, these advancements establish Co-DINO as the state of the art in object detection.
In contrast to Co-DINO and other classical DETR variants, RT-DETR reduces the computational burden of self-attention. Instead of applying self-attention across all backbone features, it restricts it to high-level, semantically rich features, which are then fused with low-level convolutional features. Furthermore, RT-DETR introduces an uncertainty-minimal query selection strategy that yields high-quality object queries by jointly optimizing their classification and localization confidence. Together, these modifications enable real-time inference while preserving competitive detection accuracy.
@@ 4.1.3 Implementation details
Table 3: Model and training specifications for pig detection. Pretrained Model indicates the dataset on which the model used for weight initialization was trained. LR Schedule denotes how the given learning rate decays over time. For RT-DETR and Co-DINO, the learning rate decays by a factor of 0.1 at the specified step. LR BB scale denotes the multiplier applied to the backbone learning rate relative to the base learning rate given by LR (max/min). Time (h) denotes the training time.
Model Backbone Pretrained Model Batch Size Epochs LR Schedule LR (max/min) LR BB Scale Time (h)
YOLOX-X DarkNet53[redmon2018yolov3] COCO[lin2014microsoft] 6 115 Cosine $\frac{6}{64}\times 10^{-2}\;\big/\;\frac{30}{64}\times 10^{-4}$ 1 15.1
YOLOv8-X CSP-Darknet[wang2020cspnet] COCO 6 110 Linear $\frac{6}{64}\times 10^{-2}\;\big/\;\frac{6}{64}\times 10^{-4}$ 1 9.7
RT-DETR-R101 ResNet[he2016deep] COCO+Obj365[shao2019objects365] 16 300 Step (at 164) $1\times 10^{-4}\;\big/\;1\times 10^{-5}$ 0.01 14.5
Co-DINO Swin-L[liu2021swin] COCO+Obj365 4 100 Step (at 90) $\frac{1}{4}\times 10^{-4}\;\big/\;\frac{1}{4}\times 10^{-5}$ 0.1 18.5
All models were trained and validated according to the dataset splits defined in Sec. 2.3. The best-performing model in terms of Average Precision (see Sec. 4.3) on the validation set was then used to obtain test performance. YOLOX and YOLOv8 were trained on a single Nvidia A100 GPU using the MMYolo toolbox [mmyolo2022]. RT-DETR was also trained on a single Nvidia A100 GPU using the official GitHub implementation, while Co-DINO was trained on four A100 GPUs using the MMDetection toolbox [chen2019mmdetection]. All models use an image resolution of $960\times 960$ pixels. The basic training configurations for Co-DINO and the largest model variants of YOLOX, YOLOv8 and RT-DETR are listed in Tab. 3. The batch size, number of epochs, and learning rate schedule were determined by experimenting with multiple configurations and selecting the one that achieved the best validation performance. The specified learning rate was obtained by linearly scaling the default value from the original publications according to the chosen batch size. Other specifications such as learning rate warmup schedules or data augmentations are in line with the original implementations. The training configurations for the smaller model variants of YOLOX, YOLOv8 and RT-DETR partly differ in the learning rate, which was also adopted from the original implementations. Otherwise, they are identical.
The speed of the trained models was measured on an Nvidia A100 GPU using the inference functionality of the MMDetection toolbox for YOLOX, YOLOv8 and Co-DINO, and using the official PyTorch code for RT-DETR. Note that inference speed could be further increased through model export and optimization frameworks such as ONNX Runtime [onnxruntime]. Here we only report the speed of the official research code. We ran the models on a set of 1000 randomly selected images, repeating the process five times to mitigate variability in performance measurements. The overall median speed per image was computed from these 5000 inferences to obtain a robust assessment. Images were resized in advance to a resolution of $960\times 960$ pixels to avoid additional overhead. Data loading from disc was not considered in the speed measurements, as it can be avoided in practical applications.
@@
The model used for training is YOLOX [ge_yolox_2021] using a checkpoint trained on COCO as initialization . A constant learning rate of $3/64\times 10^{-2}$ with a batch size of 3 was used in all 14 variants to ensure comparability. The model was trained with Mosaic and Mixup data augmentations in all epochs. Otherwise, the training settings were identical to the ones that were used to train the “YOLOX-L” model in the original YOLOX work [ge_yolox_2021]. Training was terminated when the Average Precision (see Sec. 4.3) on the test set did not increase for 20 consecutive epochs. Performance metrics are reported for the checkpoint with the highest Average Precision score.
@@ 4.3 Evaluation metrics
Model performance was evaluated with the commonly used Average Precision metric [everingham_pascal_2010]. To construct this metric, a one-to-one matching between predicted and ground-truth bounding boxes is established by minimizing the IoU across all pairs. This is done using the Hungarian algorithm [kuhn1955hungarian]. The algorithm ensures that each predicted bounding box can be matched to at most one ground-truth bounding box and vice versa. The resulting matches are subsequently filtered according to a minimum required IoU of $\alpha$. This prevents matches of predictions and ground truths that have little or no overlap. Next, all $N_{\text{pred}}$ predicted bounding boxes in the dataset that is evaluated are ordered by confidence in descending order. Then, precision and recall based on the $n$ most confident predictions and a specific $\alpha$ value can be calculated as:
@@
$\text{AP}_{\alpha}=\int_{0}^{1}\text{P}_{\alpha}(r)\,dr$
Following the literature on object detection, we report two metrics in this work: $\text{AP}_{0.5}$ and Average Precision (AP). $\text{AP}_{0.5}$ is simply $\text{AP}_{\alpha}$ computed with a minimum required IoU for a match of $\alpha=0.5$. Since the required IoU is relatively low, this metric measures a model’s ability to roughly detect objects without taking into account localization accuracy. In contrast, AP is computed by averaging over multiple IoU thresholds: